# The non-existence of some Galois representations of moderate degree in small characteristic

### v1

@@ 4. The self-dual five-dimensional case (Cor B)
Assume GRH. There exist no self-dual $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{5}(\overline{{\mathbb{F}}}_{2})$ of length $2$.
@@
${\rm O}(B)(F):=\{g\in{\rm GL}_{2n+1}(F)\ |\ B(gx,gy)=B(x,y),\ \forall x,y\in F^{\oplus 2n+1}\}.$
By [BHRD, p.23, Theorem 1.5.41], ${\rm O}(B)(F)\simeq{\rm O}(J)(F)$ where $J$ is the anti-diagonal matrix whose anti-diagonal entries are all 1. By direct computation, it is easy to see that each element of ${\rm O}(J)(F)$ can be written as $\left(\begin{array}[]{ccc}A&0&B\\ 0&1&0\\ C&0&D\end{array}\right)$ with $\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\in{\rm Sp}_{2n}(F)$. This is a special feature in the case of even characteristic. Thus, ${\rm O}(B)(F)\simeq{\rm O}(J)(F)\simeq{\rm Sp}_{2n}(F)$.
@@
Since $\overline{\rho}$ is $2$-ramified, the determinant representation $\det(\overline{\rho})$ factors though pro $2$-group ${\rm Gal}(\mathbb{Q}(\zeta_{2^{\infty}})/\mathbb{Q})$. Therefore, $\det(\overline{\rho})$ has to be trivial since its image inside $\overline{{\mathbb{F}}}^{\times}_{2}$ has odd cardinality. Thus, by assumption, $\overline{\rho}^{\vee}\simeq\overline{\rho}$. It yields
$F\subset{\rm End}_{F[G_{\mathbb{Q}}]}(\overline{\rho})\simeq{\rm Bil}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho}^{\vee},F)\simeq{\rm Bil}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)=$
${\rm Sym}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)\oplus{\rm Alt}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)$
where for $F[G_{\mathbb{Q}}]$-modules $V,V_{1}$, and $V_{2}$, ${\rm Bil}_{F[G_{\mathbb{Q}}]}(V_{1}\times V_{2},F)$ stands for the space of all $F[G_{\mathbb{Q}}]$-bilinear forms on $V_{1}\times V_{2}$ over $F$ and ${\rm Sym}_{F[G_{\mathbb{Q}}]}(V\times V)$ (resp. ${\rm Alt}_{F[G_{\mathbb{Q}}]}(V\times V)$) stands for the space of all $F[G_{\mathbb{Q}}]$-symmetric (resp. -alternating) bilinear forms on $V\times V$ over $F$. Since $\overline{\rho}$ has odd dimension, ${\rm Alt}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)=0$. Thus, ${\rm Sym}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)$ contains an element which comes from $F^{\times}\subset({\rm End}_{F[G_{\mathbb{Q}}]}(\overline{\rho}))^{\times}$ and it means that there exists a non-degenerate symmetric form $B$ such that ${\rm Im}(\overline{\rho})\subset{\rm O}(B)(F)\simeq{\rm O}(J)(F)\subset{\rm GL}_{5}(F)$. By using the description of ${\rm O}(J)(F)$, there exists a mod 2 Galois representation $\overline{\rho}_{1}:G_{\mathbb{Q}}\longrightarrow{\rm Sp}_{4}(F)\subset{\rm GL}_{4}(F)$ such that $\overline{\rho}\simeq\overline{\rho}_{1}\oplus\mathbf{1}$. Since $\overline{\rho}$ is of length 2, $\overline{\rho}_{1}$ is irreducible over $F$. Thus, one can apply Theorem A to deduce the claim. ∎
5. Non-existence of ${\rm GSp}_{4}(\mathbb{F}_{2})$-valued representations (Thm C)
In this section, we show the non-existence of irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\mathbb{F}_{2})$. The method may be useful for $G_{K}$ for some other number fields $K$ and has potential applications to representations arising from the study of $2$-torsion points on abelian surfaces over $K$.
@@ 5. Non-existence of ${\rm GSp}_{4}(\mathbb{F}_{2})$-valued representations (Thm C)
$\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\mathbb{F}_{2}),$
@@
$\tau_{f}:{\rm Gal}(\mathbb{Q}_{f}/\mathbb{Q})\hookrightarrow S_{6}\simeq{\rm GSp}_{4}(\mathbb{F}_{2}).$
@@
Consider $H:={\rm Im}(\overline{\rho})$ as a subgroup of $S_{6}$ by the identification (3). Using Magma (see Appendix A) we find that the absolute irreducibility assumption restricts the order of $H$ as follows:
@@
  • $H$ is conjugate to ${\rm Aut}(\{1,2,3\})\times{\rm Aut}(\{4,5,6\})$ in $S_{6}$; then $\overline{\rho}$ is reducible as observed in [TY, Section 3.1], so we may discard this case (which does occur for the first conjugacy class with |H|=36|H|=36).∎
@@
Any irreducible $2$-ramified representation $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\mathbb{F}_{2})$ is absolutely irreducible.
@@
${\rm rad}(\overline{{\mathbb{F}}}_{2}[{\rm Im}(\overline{\rho})])={\rm rad}(\mathbb{F}_{2}[{\rm Im}(\overline{\rho})])\otimes_{\mathbb{F}_{2}}\overline{{\mathbb{F}}}_{2},$
and ${\rm rad}(\mathbb{F}_{2}[{\rm Im}(\overline{\rho})])=0$ since $\overline{\rho}$ is irreducible over $\mathbb{F}_{2}$.
@@
Let $\overline{\rho}_{1}:G_{\mathbb{Q}}\longrightarrow\overline{{\mathbb{F}}}^{\times}_{2}$ be a one-dimensional component. Since it is unramified outside $2$, it factors through the profinite $2$-group ${\rm Gal}(\mathbb{Q}(\zeta_{2^{\infty}})/\mathbb{Q})$. On the other hand, its image lies in $\mathbb{F}^{\times}_{2^{n}}$ for some $n\in\mathbb{N}$, so has odd order. Therefore $\overline{\rho}_{1}$ must be trivial.
@@
There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\mathbb{F}_{2})$.
@@
By Proposition 5.3, we may assume that $\overline{\rho}$ is absolutely irreducible. We can therefore apply Proposition 5.2 and get an irreducible polynomial $f(x)$ over $\mathbb{Q}$ of degree 55 or 66 such that $\tau_{f}\simeq\overline{\rho}$. Then $\mathbb{Q}_{f}$ is a quintic of sextic 22-ramified number field, which is ruled out by Proposition 2.1. ∎
6. Non-existence of ${\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$-valued representations (Thm D, Cor E)
We now tackle the more general case of representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$.
@@ 6. Non-existence of ${\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$-valued representations (Thm D, Cor E)
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ be an irreducible $2$-ramified representation. Then ${\rm Im}(\overline{\rho})$ is either isomorphic to the Suzuki group ${\rm Sz}(\mathbb{F}_{2^{r}})$ for some odd $r>1$, or is equal to ${\rm Sp}_{4}(\mathbb{F}_{2^{s}})$ for some $s>1$.
In particular, there exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ such that $|{\rm Im}(\overline{\rho})|<29120$.
The following result is well-known:
@@
Let $L/\mathbb{Q}$ be a quadratic extension whose class number $|{\rm Cl}(L)|$ is a power of $2$. If $2$ is not inert in $O_{L}$, then any finite abelian extension of $L$ unramified outside $2$ has degree $2^{n}$ for some integer $n\geqslant 0$. In particular, there exist no non-trivial $2$-ramified characters $\overline{\rho}:G_{L}\longrightarrow\overline{{\mathbb{F}}}^{\times}_{2}$.
@@
Let $K/L$ be a finite abelian $2$-ramified extension. By class field theory, there exists an ideal $I\subset O_{L}$ dividing a power of the ideal $2O_{L}$ such that $K$ is the ring class field with modulus $I$. Letting $H$ denote the Hilbert class field of $L$, we have the natural surjection ${\rm Gal}(K/L)\longrightarrow{\rm Gal}(H/L)$. Its kernel is described by class field theory ([KKS, p.106, Proposition 6.114]) as
${\rm Coker}\Big{(}\mathcal{O}^{\times}_{L}\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}\Big{(}\bigoplus_{v:\text{real place}}\mathbb{R}^{\times}/\mathbb{R}^{\times}_{>0}\Big{)}\oplus(\mathcal{O}_{L}/I)^{\times}\Big{)};$
here $\alpha$ is the diagonal map whose components are induced from the natural maps $\mathcal{O}^{\times}_{L}\longrightarrow(\mathcal{O}_{L}/I)^{\times}$ and $\mathcal{O}^{\times}_{L}\longrightarrow L^{\times}_{v}=\mathbb{R}^{\times}$ if $v$ is a real place while we ignore the complex places. By assumption, the orders of both this group and the class group ${\rm Gal}(H/L)\simeq{\rm Cl}(L)$ are powers of $2$. Hence $[K:L]$ is a power of $2$.
For the second claim, notice that ${\rm Im}(\overline{\rho})$ has odd order. Thus, the claim is obvious from the previous result. ∎
@@
Before we tackle Theorem D, we recall the classification of the maximal subgroups $H$ inside $G_{r}:={\rm Sp}_{4}(\mathbb{F}_{2^{r}}),\ r>1$ (see [DZ, p.648] or [BHRD, Section 7.2]):
@@
    $H$ is reducible (over $\overline{{\mathbb{F}}}_{2}$), when viewed as a representation with values in ${\rm GL}_{4}(\overline{{\mathbb{F}}}_{2})$ via the natural inclusion $H\subset G_{r}\subset{\rm GL}_{4}(\overline{{\mathbb{F}}}_{2})$;
@@
    $H={\rm SO}^{+}(4,\mathbb{F}_{2^{r}}):=\{g\in{\rm SL}_{4}(\mathbb{F}_{2^{r}})\ |\ Q(gx)=Q(x)\}$ where $Q:\mathbb{F}^{\oplus 4}_{2^{r}}\longrightarrow\mathbb{F}_{2^{r}}$ is the quadratic map defined by $Q(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{4}+x_{2}x_{3}$;
@@
    $H={\rm SO}^{-}(4,\mathbb{F}_{2^{r}}):=\{g\in{\rm SL}_{4}(\mathbb{F}_{2^{r}})\ |\ Q(gx)=Q(x)\}$ where $Q:\mathbb{F}^{\oplus 4}_{2^{r}}\longrightarrow\mathbb{F}_{2^{r}}$ is the quadratic map defined by $Q(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{4}+x_{2}^{2}+\mu x_{3}^{2}$ for any $\mu\in\mathbb{F}_{2^{r}}$ such that $x^{2}+x+\mu$ is irreducible over $\mathbb{F}_{2^{r}}$;
@@
    $H={\rm Sz}(\mathbb{F}_{2^{r}})$, a Suzuki group (occurs only when $r$ is odd);
@@
$|{\rm SO}^{+}(\mathbb{F}_{2^{r}})|$ $=2^{2r+1}(2^{r}-1)^{2},$
$|{\rm SO}^{-}(\mathbb{F}_{2^{r}})|$ $=2^{2r+1}(2^{r}-1)(2^{r}+1),$
$|{\rm Sz}(\mathbb{F}_{2^{r}})|$ $=2^{2r}(2^{r}-1)(2^{2r}+1).$
In particular, the orthogonal groups ${\rm SO}^{\pm}(\mathbb{F}_{2^{r}})$ are proper subgroups of $G_{r}$.
The maximal subgroups of the Suzuki group ${\rm Sz}(\mathbb{F}_{2^{r}})\subset G_{r}\subset{\rm GL}_{4}(\mathbb{F}_{2^{r}})$ with $r=2s+1>1$ odd are classified in [BHRD, p.385, Table 8.16] as follows:
@@
    ${\rm Sz}(\mathbb{F}_{2^{r}})\cap B$ where $B$ is the upper triangular Borel subgroup of ${\rm Sp}_{4}$;
@@
    ${\rm Sz}(\mathbb{F}_{2^{r^{\prime}}})$ with $r^{\prime}|r$.
@@
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ be an irreducible Galois representation unramified outside $2$. Put $G={\rm Im}(\overline{\rho})$. As in the proof of Proposition 5.3, we see that the similitude character of $\overline{\rho}$ is trivial. Hence $G\subset{\rm Sp}_{4}(\overline{{\mathbb{F}}}_{2})$. Further, by Theorem C, we may assume $G\subset G_{r}$ for some minimal $r>1$.

### v2

@@ 4. The self-dual five-dimensional case (Cor B)
Assume GRH. There exist no self-dual $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{5}(\overline{\mathbb{F}}_{2})$ of length $2$.
@@
$\mathrm{O}(B)(F):=\{g\in\mathrm{GL}_{2n+1}(F)\ |\ B(gx,gy)=B(x,y),\ \forall x,y\in F^{\oplus 2n+1}\}.$
By [BHRD, p.23, Theorem 1.5.41], $\mathrm{O}(B)(F)\simeq\mathrm{O}(J)(F)$ where $J$ is the anti-diagonal matrix whose anti-diagonal entries are all 1. By direct computation, it is easy to see that each element of $\mathrm{O}(J)(F)$ can be written as $\left(\begin{array}[]{ccc}A&0&B\\ 0&1&0\\ C&0&D\end{array}\right)$ with $\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\in\mathrm{Sp}_{2n}(F)$. This is a special feature in the case of even characteristic. Thus, $\mathrm{O}(B)(F)\simeq\mathrm{O}(J)(F)\simeq\mathrm{Sp}_{2n}(F)$.
@@
Since $\overline{\rho}$ is $2$-ramified, the determinant representation $\det(\overline{\rho})$ factors though pro-$2$-group $\mathrm{Gal}(\mathbb{Q}(\zeta_{2^{\infty}})/\mathbb{Q})$. Therefore, $\det(\overline{\rho})$ has to be trivial since its image inside $\overline{\mathbb{F}}^{\times}_{2}$ has odd cardinality. Thus, by assumption, $\overline{\rho}^{\vee}\simeq\overline{\rho}$. It yields
$F\subset\mathrm{End}_{F[G_{\mathbb{Q}}]}(\overline{\rho})\simeq\mathrm{Bil}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho}^{\vee},F)\simeq\mathrm{Bil}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)=\\ \mathrm{Sym}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)\oplus\mathrm{Alt}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F),$
where $\mathrm{Bil}$, $\mathrm{Sym}$, resp. $\mathrm{Alt}$ denote the spaces of all $F[G_{\mathbb{Q}}]$-bilinear, symmetric bilinear, resp. alternating bilinear forms on the corresponding $F[G_{\mathbb{Q}}]$-modules. Since $\overline{\rho}$ has odd dimension, $\mathrm{Alt}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)=0$. Thus, $\mathrm{Sym}_{F[G_{\mathbb{Q}}]}(\overline{\rho}\times\overline{\rho},F)$ contains an element which comes from $F^{\times}\subset(\mathrm{End}_{F[G_{\mathbb{Q}}]}(\overline{\rho}))^{\times}$ and it means that there exists a non-degenerate symmetric form $B$ such that
$\mathrm{Im}(\overline{\rho})\subset\mathrm{O}(B)(F)\simeq\mathrm{O}(J)(F)\subset\mathrm{GL}_{5}(F).$
By using the description of $\mathrm{O}(J)(F)$, there exists a mod 2 Galois representation $\overline{\rho}_{1}:G_{\mathbb{Q}}\longrightarrow\mathrm{Sp}_{4}(F)\subset\mathrm{GL}_{4}(F)$ such that $\overline{\rho}\simeq\overline{\rho}_{1}\oplus\mathbf{1}$. Since $\overline{\rho}$ is of length 2, $\overline{\rho}_{1}$ is irreducible over $F$. Thus, one can apply Theorem A to deduce the claim. ∎
5. Non-existence of $\mathrm{GSp}_{4}(\mathbb{F}_{2})$-valued representations (Thm C)
In this section, we show the non-existence of irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{2})$. The method may be useful for $G_{K}$ for some other number fields $K$ and has potential applications to representations arising from the study of $2$-torsion points on abelian surfaces over $K$.
@@ 5. Non-existence of $\mathrm{GSp}_{4}(\mathbb{F}_{2})$-valued representations (Thm C)
$\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{2}),$
@@
$\tau_{f}:\mathrm{Gal}(\mathbb{Q}_{f}/\mathbb{Q})\hookrightarrow S_{6}\simeq\mathrm{GSp}_{4}(\mathbb{F}_{2}).$
@@
Consider $H:=\mathrm{Im}(\overline{\rho})$ as a subgroup of $S_{6}$ by the identification (3). Using Magma (see Appendix A) we find that the absolute irreducibility assumption restricts the order of $H$ as follows:
@@
  • $H$ is conjugate to $\mathrm{Aut}(\{1,2,3\})\times\mathrm{Aut}(\{4,5,6\})$ in $S_{6}$; then $\overline{\rho}$ is reducible as observed in [TY, Section 3.1], so we may discard this case (which does occur for the first conjugacy class with |H|=36|H|=36).∎
@@
Any irreducible $2$-ramified representation $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{2})$ is absolutely irreducible.
@@
$\mathrm{rad}(\overline{\mathbb{F}}_{2}[\mathrm{Im}(\overline{\rho})])=\mathrm{rad}(\mathbb{F}_{2}[\mathrm{Im}(\overline{\rho})])\otimes_{\mathbb{F}_{2}}\overline{\mathbb{F}}_{2},$
and $\mathrm{rad}(\mathbb{F}_{2}[\mathrm{Im}(\overline{\rho})])=0$ since $\overline{\rho}$ is irreducible over $\mathbb{F}_{2}$.
@@
Let $\overline{\rho}_{1}:G_{\mathbb{Q}}\longrightarrow\overline{\mathbb{F}}^{\times}_{2}$ be a one-dimensional component. Since it is unramified outside $2$, it factors through the profinite $2$-group $\mathrm{Gal}(\mathbb{Q}(\zeta_{2^{\infty}})/\mathbb{Q})$. On the other hand, its image lies in $\mathbb{F}^{\times}_{2^{n}}$ for some $n\in\mathbb{N}$, so has odd order. Therefore $\overline{\rho}_{1}$ must be trivial.
@@
There exist no representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{2})$ that are irreducible and $2$-ramified.
@@
By Proposition 5.3, we may assume that $\overline{\rho}$ is absolutely irreducible. We can therefore apply Proposition 5.2 and get an irreducible polynomial $f(x)$ over $\mathbb{Q}$ of degree 55 or 66 such that $\tau_{f}\simeq\overline{\rho}$. Then $\mathbb{Q}_{f}$ is a quintic or sextic 22-ramified number field, which is ruled out by Proposition 2.1. ∎
6. Non-existence of $\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$-valued representations (Thm D, Cor E)
We now tackle the more general case of representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$.
@@ 6. Non-existence of $\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$-valued representations (Thm D, Cor E)
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ be an irreducible $2$-ramified representation. Then $\mathrm{Im}(\overline{\rho})$ is either isomorphic to the Suzuki group $\mathrm{Sz}(\mathbb{F}_{2^{r}})$ for some odd $r>1$, or is equal to $\mathrm{Sp}_{4}(\mathbb{F}_{2^{s}})$ for some $s>1$.
Hence there exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ such that $|\mathrm{Im}(\overline{\rho})|<29120$.
We include the following well-known result for completeness:
@@
Let $L/\mathbb{Q}$ be a quadratic extension whose class number $|\mathrm{Cl}(L)|$ is a power of $2$. If $2$ is not inert in $O_{L}$, then any finite abelian extension of $L$ unramified outside $2$ has degree $2^{n}$ for some integer $n\geqslant 0$. In particular, there exist no non-trivial $2$-ramified characters $\overline{\rho}:G_{L}\longrightarrow\overline{\mathbb{F}}^{\times}_{2}$.
@@
Let $K/L$ be a finite abelian $2$-ramified extension. By class field theory, there exists an ideal $I\subset O_{L}$ dividing a power of the ideal $2O_{L}$ such that $K$ is the ring class field with modulus $I$. Letting $H$ denote the Hilbert class field of $L$, we have the natural surjection $\mathrm{Gal}(K/L)\longrightarrow\mathrm{Gal}(H/L)$. Its kernel is described by class field theory ([KKS, p.106, Proposition 6.114]) as
$\mathrm{coker}\Big(\mathcal{O}^{\times}_{L}\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}\Big(\bigoplus_{v:\text{real place}}\mathbb{R}^{\times}/\mathbb{R}^{\times}_{>0}\Big)\oplus(\mathcal{O}_{L}/I)^{\times}\Big);$
here $\alpha$ is the diagonal map whose components are induced from the natural maps $\mathcal{O}^{\times}_{L}\longrightarrow(\mathcal{O}_{L}/I)^{\times}$ and $\mathcal{O}^{\times}_{L}\longrightarrow L^{\times}_{v}=\mathbb{R}^{\times}$ if $v$ is a real place while we ignore the complex places. By assumption, the orders of both this group and the class group $\mathrm{Gal}(H/L)\simeq\mathrm{Cl}(L)$ are powers of $2$. Hence $[K:L]$ is a power of $2$.
For the second claim, notice that $\mathrm{Im}(\overline{\rho})$ has odd order. Thus, the claim is obvious from the previous result. ∎
@@
Before we tackle Theorem D, we recall the classification of the maximal subgroups $H$ inside $G_{r}:=\mathrm{Sp}_{4}(\mathbb{F}_{2^{r}}),\ r>1$ (see [DZ, p.648] or [BHRD, Section 7.2]):
@@
    $H$ is reducible (over $\overline{\mathbb{F}}_{2}$), when viewed as a representation with values in $\mathrm{GL}_{4}(\overline{\mathbb{F}}_{2})$ via the natural inclusion $H\subset G_{r}\subset\mathrm{GL}_{4}(\overline{\mathbb{F}}_{2})$;
@@
    $H=\mathrm{SO}^{+}(4,\mathbb{F}_{2^{r}}):=\{g\in\mathrm{SL}_{4}(\mathbb{F}_{2^{r}})\ |\ Q(gx)=Q(x)\}$ where $Q:\mathbb{F}^{\oplus 4}_{2^{r}}\longrightarrow\mathbb{F}_{2^{r}}$ is the quadratic map defined by $Q(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{4}+x_{2}x_{3}$;
@@
    $H=\mathrm{SO}^{-}(4,\mathbb{F}_{2^{r}}):=\{g\in\mathrm{SL}_{4}(\mathbb{F}_{2^{r}})\ |\ Q(gx)=Q(x)\}$ where $Q:\mathbb{F}^{\oplus 4}_{2^{r}}\longrightarrow\mathbb{F}_{2^{r}}$ is the quadratic map defined by $Q(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{4}+x_{2}^{2}+\mu x_{3}^{2}$ for any $\mu\in\mathbb{F}_{2^{r}}$ such that $x^{2}+x+\mu$ is irreducible over $\mathbb{F}_{2^{r}}$;
@@
    $H=\mathrm{Sz}(\mathbb{F}_{2^{r}})$, a Suzuki group (occurs only when $r$ is odd);
@@
$|\mathrm{SO}^{+}(\mathbb{F}_{2^{r}})|$ $=2^{2r+1}(2^{r}-1)^{2},$
$|\mathrm{SO}^{-}(\mathbb{F}_{2^{r}})|$ $=2^{2r+1}(2^{r}-1)(2^{r}+1),$
$|\mathrm{Sz}(\mathbb{F}_{2^{r}})|$ $=2^{2r}(2^{r}-1)(2^{2r}+1).$
In particular, the orthogonal groups $\mathrm{SO}^{\pm}(\mathbb{F}_{2^{r}})$ are proper subgroups of $G_{r}$.
The maximal subgroups of the Suzuki group $\mathrm{Sz}(\mathbb{F}_{2^{r}})\subset G_{r}\subset\mathrm{GL}_{4}(\mathbb{F}_{2^{r}})$ with $r=2s+1>1$ odd are classified in [BHRD, p.385, Table 8.16] as follows:
@@
    $\mathrm{Sz}(\mathbb{F}_{2^{r}})\cap B$ where $B$ is the upper triangular Borel subgroup of $\mathrm{Sp}_{4}$;
@@
    $\mathrm{Sz}(\mathbb{F}_{2^{r^{\prime}}})$ with $r^{\prime}|r$.
@@
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ be an irreducible Galois representation unramified outside $2$. Put $G=\mathrm{Im}(\overline{\rho})$. As in the proof of Proposition 5.3, we see that the similitude character of $\overline{\rho}$ is trivial. Hence $G\subset\mathrm{Sp}_{4}(\overline{\mathbb{F}}_{2})$. Further, by Theorem C, we may assume $G\subset G_{r}$ for some minimal $r>1$.